# Denoising as Adaptation: Noise-Space Domain Adaptation for Image Restoration

### v2

@@ 2 Related Work
Image Restoration. Image restoration aims to recover images degraded by factors like noise, blur, or data loss. Driven largely by the capabilities of various networks (Dong et al., 2014; 2015; Zamir et al., 2022; Liang et al., 2021), significant advancements have been made in sub-fields such as image denoising (Zhang et al., 2021; Ren et al., 2021; Guo et al., 2019; Kim et al., 2020), image deblurring (Kupyn et al., 2018; Suin et al., 2020; Zhang et al., 2019), and image deraining (Jiang et al., 2020; Purohit et al., 2021; Ren et al., 2019; Yang et al., 2017). In image restoration, loss functions are essential for training models. For example, the $L1$ loss minimizes average absolute pixel differences, ensuring pixel-wise accuracy. Perceptual loss uses pre-trained networks to compare high-level features, ensuring perceptual similarity. Adversarial loss involves a discriminator distinguishing between real and synthetic images, pushing the generator to create more realistic outputs. However, restoration models trained on synthetic images with these conventional loss functions still cannot escape from a significant drop in performance when applied to real-world domains.
@@
Diffusion Model. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021) have gained significant attention in generative modeling. They work by gradually transforming a simple distribution into a complex distribution in a series of steps, reversing the diffusion process. This approach shows remarkable success in text-to-image generation (Saharia et al., 2022b; Ruiz et al., 2023) and image restoration (Saharia et al., 2022a; c; Yue et al., 2023). Often, conditions are fed to the diffusion model for conditional generation, such as text (Rombach et al., 2021), class label (Ho & Salimans, 2022), visual prompt (Bar et al., 2022), and low-resolution image (Wang et al., 2023), to facilitate the approximation of the target distribution. Some recent works propose to adapt diffusion models for image restoration and its related tasks such as blind JPEG restoration (Welker et al., 2024), open-set image restoration (Gou et al., 2024), and classification of degraded images (Daultani et al., 2024). However, they require the diffusion model in both the training and inference stages. In this work, we show that the diffusion’s forward denoising process has the potential to serve as a training proxy task to improve the generalization ability of the image restoration model.
@@ 3 Methodology
Image Restoration Baseline. The image restoration network can be generally formulated as a deep neural network $G(\cdot;{\bm{\theta}}_{G})$ with learnable parameter $\bm{\theta}_{G}$. This network is trained to predict the ground truth image ${\bm{y}}^{s}$ from its degraded observation $\bm{x}^{s}$ on the synthetic domain. The proposed noise space domain adaptation is not limited to a specific type of network architecture. One can choose from existing networks such as DnCNN (Zhang et al., 2017), U-Net (Yue et al., 2019), RCAN (Zhang et al., 2018c), and SwinIR (Liang et al., 2021). The approach is also orthogonal to existing loss functions used in image restoration, e.g., $L_{1}$ or $L_{2}$ loss, Charbonnier loss (Zamir et al., 2021), perceptual loss (Johnson et al., 2016; Zhang et al., 2018b), and adversarial loss (Wang et al., 2018; Kupyn et al., 2018). To better validate the generality of the proposed approach, we adopt the widely used U-Net architecture and the Charbonnier loss, denoted as $\mathcal{L}_{Res}$, as our baseline. In the joint training, the diffusion model is trained using a diffusion objective, $\mathcal{L}_{Dif}$, while the restoration network is updated using both the $\mathcal{L}_{Res}$ and $\mathcal{L}_{Dif}$. The diffusion model is discarded after training.
@@ 3.1 Noise-Space Domain Adaptation
where $\tilde{\bm{y}}^{s}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{s}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, $\bm{\epsilon}\sim N(0,\bm{I})$, $\bar{\alpha}_{t}$ is the hyper-parameter of the noise schedule, and $\mathbf{C}(\cdot,\cdot)$ denotes the concatenation operation along the channel dimension. During the joint training, supervision from the diffusion loss in Eq. 1 will back-propagate to the conditions $\hat{\bm{y}}^{s}$ and $\hat{\bm{y}}^{r}$ if they are under-restored, i.e., far away from the expected distribution. This encourages the preceding restoration network to align $\hat{\bm{y}}^{s}$ and $\hat{\bm{y}}^{r}$ as closely as possible to the target clean distribution.
@@ 3.2 Eliminating Shortcut Learning in Diffusion
where $\delta$ denotes a predefined margin to separate the positive and negative samples. In this way, the loss of diffusion model takes the mean of Eq. 1 and Eq. 3. By implementing the above strategies, we challenge the diffusion model to distinguish between synthetic and real conditions based on trivial solutions, thereby encouraging both to align with the target clean distribution.
In the above formulation, the synthetic restored image of the condition, denoted as $\hat{\bm{y}}^{s}$, and the input to the diffusion model, represented as $\tilde{\bm{y}}^{s}$, form a pair of data with evident pixel-wise similarity. This similarity can potentially mislead the diffusion model to ignore the real restored image $\hat{\bm{y}}^{r}$ in condition as analyzed in Fig. 2. It is important to note that the target distribution encapsulates the domain knowledge of high-quality clean images, including but not limited to the ground truth images in the synthetic dataset. Motivated by this observation, the proposed method can be further extended by replacing the noisy input $\tilde{\bm{y}}^{s}$ with $\tilde{\bm{y}}^{c}$, defined as $\tilde{\bm{y}}^{c}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{c}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, where ${\bm{y}}^{c}$ is randomly sampled from an unpaired extensive high-quality image dataset. This strategy disrupts the pixel-wise similarity between the synthetic condition and the diffusion input, thus enforcing the diffusion model to guide both the synthetic and real conditions predicted by the restoration network at the domain level. We will provide an ablation on this setting in Appendix A4.1.
@@ 3.3 Training
In the proposed training strategy, the restoration and diffusion models are jointly optimized by:
$\mathcal{L}=\mathcal{L}_{Res}+\lambda_{Dif}\left[\frac{\mathcal{L}_{Dif}+\mathcal{L}_{Con}}{2}\right].$ (4)
@@ 4 Experiments
Figure 4: Visual comparison of the image denoising task on SIDD test dataset (Abdelhamed et al., 2018).
Table 1: Quantitative evaluation of the image denoising task on SIDD test dataset (Abdelhamed et al., 2018). syn, real, both denote the model is trained on synthetic, real-world (w/o GT), and both synthetic and real-world (w/o GT) datasets, respectively. The best score is highlighted.
@@ 4.1 Comparisons with State-of-the-Art Methods
Figure 5: Visual comparison of the image deraining and image deblurring tasks on SPA Wang et al. (2019) and RealBlur-J Rim et al. (2020) test datasets.
@@
Analysis. From the above results, we can observe that the proposed method enables noticeable improvements beyond the Vanilla baseline (trained only with synthetic datasets) on the tasks involved with high-frequency noises, such as image denoising. In particular, $+8.13/0.3070$ improvements on PSNR/SSIM metrics are achieved. We argue that the target of image denoising naturally fits that of the forward denoising process in the diffusion model. It is more sensitive to other Gaussian-like noises with respect to the pre-sampled noise space. Thus, an intense diffusion loss would be back-propagated if the conditioned images are under-restored, and the preceding restoration network tries to eliminate the noises on both the synthetic and real-world images as much as possible.
@@ 4.2 Ablation Studies
To evaluate the effectiveness of different components in the proposed method, we conduct ablation studies regarding the sampled noise levels of the diffusion model, determined by the time-step $t$, and the training strategies to avoid shortcut learning, as shown in Tab. 4 and Fig. 6. Concretely, with low noise intensity, e.g., $t\in[1,100]$, it is easy for the diffusion model to discriminate the similarity of paired synthetic data even when the restored conditions are under-restored. As a result, the shortcut learning occurs earlier during the training process and the real-world degraded image is heavily corrupted by the restoration network, of which most all details are filtered. On the other hand, when the intensity of the sampled noise is high, e.g., $t\in[900,1000]$, the diffusion model is hard to converge and the whole framework has fallen into a local optimum. By sampling the noise from a more diverse range with $t\in[1,1000]$, the restored results can be gradually adapted to the target clean distribution. Moreover, the generalization ability of the restoration network gains further improvement using the designed channel shuffling layer (CS) and residual-swapping contrastive learning strategy (RS), which effectively eliminates the shortcut learning of the diffusion model. Therefore, higher restoration performance on real-world images and more realistic visual appearance can be observed from (d) to (e) and (f) in Tab. 4 and Fig. 6. We also demonstrate that both synthetic data and real data are indispensable for domain adaptation in diffusion, excluding each of them would lead to dramatic degradation in real-world performance (shown in the last two rows in Tab. 4). Particularly for excluding the real data, the performance is almost degraded to that of the Vanilla model. More analysis can be found in the Appendix A3.
@@ 4.3 Scalability
Comparisons. In this work, we aim to present a general domain adaptation strategy for various image restoration tasks, which is scalable to any restoration network. In particular, a basic and lightweight U-Net is used to validate the effectiveness of the proposed strategy. However, such an architecture essentially limits the upper bound of the restoration performance compared to some recent promising self-supervised works (Jang et al., 2021; Lee et al., 2022; Jang et al., 2024) tailored to specific tasks.
@@
Discussion. Compared to the self-supervised methods (Chen et al., 2023; Ren et al., 2020; Huo et al., 2023; Jang et al., 2021; Lee et al., 2022), our work shows the following unique strengths: it is not bounded to the specific tasks; it is free to the prior knowledge of underlying noise distribution and degradation mode; and it is friendly to the type of preceding restoration networks.
We also argue the difference between domain adaptation and self-supervised learning methods: Domain adaptation transfers knowledge from one domain to another with different data distributions, improving performance in new, unseen environments. Self-supervised learning, on the other hand, learns from unlabeled data by generating pseudo-labels or exploring the target distribution from the data itself. Both approaches reduce the reliance on large amounts of labeled data but address different challenges: domain adaptation focuses on bridging domain gaps, and self-supervised learning leverages data’s inherent structure.
Peformance vs. Complexity. We then validate the scalability of the proposed method using different variants of U-Net-based image restoration networks and other types of architectures, such as the Transformer-based network (Wang et al., 2022). In particular, we classify these networks based on their model sizes and obtain: Unet-T, Unet-S (the model applied in Sec. 4.1), Unet-B, Uformer-T, Uformer-S, and Uformer-B. More network details are listed in the Appendix A1.3. The quantitative results of PSNR vs. computational costs are shown in Fig. 7. As we can observe, as the complexity and parameter increase, the vanilla restoration network (orange elements) tends to overfit the training synthetic dataset and perform worse on the test real-world dataset. In contrast, the proposed method can improve the generalization ability of restoration models with various sizes (blue elements). It is also interesting that for each type of architecture, our method can facilitate better performance as the complexity of the restoration network increases, demonstrating its effectiveness in addressing the overfitting problem of large models.
@@ 5 Conclusion
. Task Ours Ours-Ex Denoising 34.71/0.9202/0.0903 33.44/0.8938/0.1064 Deraining 34.39/0.9571/0.0462 34.20/0.9587/0.0444 Deblurring 26.46/0.8048/0.1363 26.44/0.8030/0.1313
@@
Figure A6: Visual results of the proposed method on DND real-world denoising test dataset (Plotz & Roth, 2017).
@@
Figure A7: Visual results of the proposed method on DND real-world denoising test dataset (Plotz & Roth, 2017).
@@
Figure A8: Visual results of the proposed method on ‘Real-Internet’ real-world deraining test dataset (Yang et al., 2017).

### v3

@@ 2 Related Work
Image Restoration aims to recover images degraded by factors like noise, blur, or data loss. Driven largely by the capabilities of various networks (Dong et al., 2014; 2015; Zamir et al., 2022; Liang et al., 2021), significant advancements have been made in sub-fields such as image denoising (Zhang et al., 2021; Ren et al., 2021; Guo et al., 2019; Kim et al., 2020; 2024; Fu et al., 2023; Kousha et al., 2022), image deblurring (Kupyn et al., 2018; Suin et al., 2020; Zhang et al., 2019), and image deraining (Jiang et al., 2020; Purohit et al., 2021; Ren et al., 2019; Yang et al., 2017). In image restoration, loss functions are essential for training models. For example, the $L1$ loss minimizes average absolute pixel differences, ensuring pixel-wise accuracy. Perceptual loss uses pre-trained networks to compare high-level features, ensuring perceptual similarity. Adversarial loss involves a discriminator distinguishing between real and synthetic images, pushing the generator to create more realistic outputs. However, the models trained on synthetic images with these conventional losses still cannot escape from a significant drop in performance when applied to real-world domains.
@@
Diffusion Model. Diffusion models (Sohl-Dickstein et al., 2015; Ho et al., 2020; Nichol & Dhariwal, 2021) have gained significant attention in generative modeling. They work by gradually transforming a simple distribution into a complex distribution in a series of steps, reversing the diffusion process. This approach shows remarkable success in text-to-image generation (Saharia et al., 2022b; Ruiz et al., 2023) and image restoration (Saharia et al., 2022a; c; Yue et al., 2023). Often, conditions are fed to the diffusion model for conditional generation, such as text (Rombach et al., 2022), class label (Ho & Salimans, 2022), visual prompt (Bar et al., 2022), and low-resolution image (Wang et al., 2024), to facilitate the approximation of the target distribution. Some recent works propose to adapt diffusion models for image restoration and its related tasks such as blind JPEG restoration (Welker et al., 2024), open-set image restoration (Gou et al., 2024), and classification of degraded images (Daultani et al., 2024). However, they require the diffusion model in both the training and inference stages. In this work, we show that the diffusion’s forward denoising process has the potential to serve as a training proxy task to improve the generalization ability of the image restoration model.
@@ 3 Methodology
Image Restoration Baseline. The restoration network can be formulated as a deep neural network $G(\cdot;{\bm{\theta}}_{G})$ with learnable parameter $\bm{\theta}_{G}$. This network is trained to predict the ground truth ${\bm{y}}^{s}$ from its degraded observation $\bm{x}^{s}$ on the synthetic domain. Our domain adaptation is not limited to a specific type of network architecture. One can choose from existing networks such as DnCNN (Zhang et al., 2017), U-Net (Yue et al., 2019), RCAN (Zhang et al., 2018b), and SwinIR (Liang et al., 2021). The approach is also orthogonal to existing loss functions used in image restoration, e.g., $L_{1}$ or $L_{2}$ loss, Charbonnier loss (Zamir et al., 2021), perceptual loss (Johnson et al., 2016), and adversarial loss (Wang et al., 2018). To better validate the generality of the proposed approach, we adopt the widely used U-Net architecture $\bm{f}_{\theta}(\cdot)$ and the Charbonnier loss $\mathcal{L}_{Res}$, as our baseline. In the joint training, the diffusion model is trained using a diffusion objective, $\mathcal{L}_{Dif}$, while the restoration network is updated using both the $\mathcal{L}_{Res}$ and $\mathcal{L}_{Dif}$. The diffusion model is discarded after training.
@@ 3.1 Noise-Space Domain Adaptation
where $\tilde{\bm{y}}^{s}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{s}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, $\bm{\epsilon}\sim N(0,\bm{I})$, $\bar{\alpha}_{t}$ is the hyper-parameter of the noise schedule, and $\mathbf{C}(\cdot,\cdot)$ denotes the channel-wise concatenation. During the joint training, supervision from the diffusion loss will back-propagate to the conditions $\hat{\bm{y}}^{s}$ and $\hat{\bm{y}}^{r}$ if they are under-restored, i.e., far away from the expected distribution. This encourages the restoration network to align $\hat{\bm{y}}^{s}$ and $\hat{\bm{y}}^{r}$ as closely as possible to the target domain. Particularly, the knowledge “leaked” by the diffusion’s input plays an important role, potentially offering degradation-free guidance to help adapt the degraded real-world images into the clean distribution. More discussions are presented in Section A5.2.
@@ 3.2 Eliminating Shortcut Learning in Diffusion
where $\delta$ denotes a predefined margin to separate the positive and negative samples. In this way, the loss of diffusion model takes the mean of Eq. 1 and Eq. 3, forming the final diffusion loss $\mathcal{L}_{Dif}$. Using the above strategies, we challenge the diffusion model to distinguish between synthetic and real conditions based on trivial solutions, encouraging both to align with the target clean distribution.
In the above formulation, the restored synthetic image of the condition, denoted as $\hat{\bm{y}}^{s}$, and the input to the diffusion model, represented as $\tilde{\bm{y}}^{s}$, form a pair of data with evident pixel-wise similarity. This similarity can potentially mislead the diffusion model to ignore the real restored image $\hat{\bm{y}}^{r}$ in condition as analyzed in Fig. 2. It is important to note that the target distribution encapsulates the domain knowledge of high-quality clean images, including but not limited to the ground truth images in the synthetic dataset. Motivated by this observation, the proposed method can be further extended by replacing the noisy input $\tilde{\bm{y}}^{s}$ with $\tilde{\bm{y}}^{c}$, defined as $\tilde{\bm{y}}^{c}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{c}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, where ${\bm{y}}^{c}$ is randomly sampled from an unpaired extensive high-quality image dataset. This strategy disrupts the pixel-wise similarity between the synthetic condition and the diffusion input, thus enforcing the diffusion model to guide both the synthetic and real conditions predicted by the restoration network at the domain level. We will provide an ablation on this setting in Appendix A4.1.
@@ 3.3 Training
In the proposed training strategy, the image restoration model is jointly optimized by:
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\mathcal{L}=\mathcal{L}_{Res}+\lambda_{Dif}\mathcal{L}_{Dif}.}}$ (4)
@@ 4 Experiments
Figure 4: Visual comparison of the image denoising task on SIDD test dataset (Abdelhamed et al., 2018). PSNR (dB) is marked for each comparison sample.
Table 1: Quantitative evaluation of the image denoising task on SIDD test dataset. syn, real, both denote the model is trained on synthetic, real-world (w/o GT), and both synthetic and real-world (w/o GT) datasets, respectively. $\mathcal{L}_{Res}$, $\mathcal{L}_{Gan}$, $\mathcal{L}_{Ori}$, and $\mathcal{L}_{Dif}$ denotes the pixel-wise restoration loss (i.e., Charbonnier loss), generative adversarial loss, original loss exploited in the paper, and the proposed diffusion loss, respectively.
@@
Train Loss $\mathcal{L}_{Res}$ $\mathcal{L}_{Res}$+$\mathcal{L}_{Gan}$ $\mathcal{L}_{Res}$+$\mathcal{L}_{Gan}$ $\mathcal{L}_{Res}$+$\mathcal{L}_{Gan}$ $\mathcal{L}_{Res}$+$\mathcal{L}_{Gan}$ $\mathcal{L}_{Ori}$ $\mathcal{L}_{Ori}$ $\mathcal{L}_{Res}$+$\mathcal{L}_{Dif}$
@@ 4.1 Comparisons with State-of-the-Art Methods
Figure 5: Visual comparison of the image deraining and image deblurring tasks on SPA Wang et al. (2019) and RealBlur-J Rim et al. (2020) test datasets. PSNR (dB) is marked for each comparison sample.
@@
Analysis. From the above results, we can observe that the proposed method enables noticeable improvements beyond the baseline on the tasks involved with high-frequency noises, such as image denoising. In particular, $+8.13/0.3070$ gains on PSNR/SSIM are achieved. We argue that the target of image denoising naturally fits that of the forward denoising process in the diffusion model. It is more sensitive to other Gaussian-like noises in the pre-sampled noise space. Thus, an intense diffusion loss would be back-propagated if the conditioned images are under-restored, and the restoration network tries to eliminate the noises on both the synthetic and real-world images as much as possible.
@@ 4.2 Ablation Studies
We conduct ablation studies regarding the sampled noise levels of the diffusion model, determined by the time-step $t$, and the training strategies to avoid shortcut learning, as shown in Tab. 4 and Fig. 6. Concretely, with low noise intensity, e.g., $t\in[1,100]$, it is easy for the diffusion model to discriminate the similarity of paired synthetic data even when the restored conditions are under-restored. As a result, the shortcut learning occurs earlier during the training process and the real-world degraded image is heavily corrupted by the restoration network, of which most details are filtered. On the other hand, when the intensity of the sampled noise is high, e.g., $t\in[900,1000]$, the diffusion model is hard to converge and the whole framework has fallen into a local optimum. By sampling the noise from a more diverse range with $t\in[1,1000]$, the restored results can be gradually adapted to the clean distribution. Moreover, the generalization ability of the restoration network gains further improvement using the designed channel shuffling layer (CS) and residual-swapping contrastive learning strategy (RS), which effectively eliminates the shortcut learning of the diffusion model. Therefore, higher restoration performance on real-world images and more realistic visual appearance can be observed from (d) to (e) and (f) in Tab. 4 and Fig. 6. We also demonstrate that both synthetic data and real data are indispensable for our domain adaptation, excluding each of them would lead to dramatic degradation in real-world performance (shown in the last two rows in Tab. 4). Particularly for excluding the real data, the performance is almost degraded to that of the Vanilla model.
@@ 4.3 Scalability
Comparisons. In this work, we aim to present a general domain adaptation strategy for various restoration tasks, which is scalable to any restoration network. In particular, a basic and lightweight U-Net is used to validate the effectiveness of our method. However, such an architecture essentially limits the upper bound of the restoration performance compared to some recent self-supervised works (Jang et al., 2021; Lee et al., 2022; Jang et al., 2024; Cai et al., 2021) tailored to specific tasks.
@@
Discussion. Compared to the self-supervised methods (Chen et al., 2023; Ren et al., 2020; Huo et al., 2023; Jang et al., 2021; Lee et al., 2022), our work shows the following unique strengths: it is not bounded to the specific tasks; it is free to the prior knowledge of underlying noise distribution and degradation mode; and it is friendly to the type of preceding restoration networks. We also argue the difference between domain adaptation and self-supervised learning methods: Domain adaptation transfers knowledge from one domain to another with different distributions, improving performance in new, unseen environments. Self-supervised learning, on the other hand, learns from unlabeled data by generating pseudo-labels or exploring the target distribution from the data itself. Both approaches reduce the reliance on large labeled data but address different challenges: domain adaptation focuses on bridging domain gaps, and self-supervised learning leverages data’s inherent structure.
Performance vs. Complexity. We validate the scalability of the proposed method using different variants of U-Net-based restoration networks and other types of architectures, such as the Transformer-based network (Wang et al., 2022). In particular, we classify these networks based on their model sizes and obtain: Unet-T, Unet-S (the model applied in Sec. 4.1), Unet-B, Uformer-T, Uformer-S, and Uformer-B. More details are listed in the Appendix. The quantitative results vs. computational costs are shown in Fig. 7. As we can observe, as the complexity increases, the vanilla restoration network (orange elements) tends to overfit the training synthetic dataset and perform worse on the test real-world dataset. In contrast, the proposed method can improve the generalization ability of restoration models with various sizes (blue elements). It is also interesting that for each type of architecture, our method can facilitate better performance as the complexity of the restoration network increases, demonstrating its effectiveness in addressing the overfitting problem of large models.
@@ 5 Conclusion
Acknowledgment
This study is supported under the RIE2020 Industry Alignment Fund – Industry Collaboration Projects (IAF-ICP) Funding Initiative, as well as cash and in-kind contribution from the industry partner(s).
@@ Acknowledgment
Task Ours Ours-Ex
Denoising 34.71/0.9202/0.0903 33.44/0.8938/0.1064
Deraining 34.39/0.9571/0.0462 34.20/0.9587/0.0444
Deblurring 26.46/0.8048/0.1363 26.44/0.8030/0.1313
@@
Figure A6: Visual results of the proposed method on unseen DND real-world denoising test dataset (Plotz & Roth, 2017).
@@
Figure A7: Visual results of the proposed method on unseen DND real-world denoising test dataset (Plotz & Roth, 2017).
@@
Figure A8: Visual results of the proposed method on unseen ‘Real-Internet’ real-world deraining test dataset (Yang et al., 2017).
A4.5 Failure Case
We show the failure case of our method and comparison methods in Fig A9. Particularly, our method fails to restore the images with challenging degraded distortions such as strong noises and out-of-distribution noises. These real-world degradations induce a significant gap compared with the synthetic dataset, burdening the learning model to effectively adapt the restored results into the clean domain.
A4.6 Analysis on Training Dynamics and Complexity
To demonstrate the impact of the introduced diffusion loss during training, we visualize the related metrics of training dynamics in Fig. A10 (left). It is easy to find that the restoration model trained only with $L1$ loss on the synthetic dataset tends to overfit quickly and performs poorly on the real-world validation set. By contrast, the diffusion loss can effectively guide the restoration model to adapt to the real-world domain in a multi-step denoising manner, consistently improving the restoration performance on the real-world validation set.
Moreover, we show the results of validating the diffusion model with different complexities in Fig. A10 (right). To be specific, we classify the diffusion models into three types based on their complexities in each layer: Diffusion-T: [32, 32, 64, 64], Diffusion-S: [32, 64, 128, 128], Diffusion-B: [64, 128, 256, 512] (exploited in this paper). As we can observe, the real-world restoration performance gains further improvements as the complexity of the diffusion model increases, i.e., from Diffusion-T to Diffusion-B. We also provide a deep analysis of diffusion loss in restoration tasks when compared with MAR (Li et al., 2024): MAR models the per-token probability distribution using a small MLP as the diffusion model. It is trained jointly with the AR model to achieve efficient image generation. In particular, the tokens are small in size and represent high-level semantic features. By contrast, our diffusion model serves for the low-level image restoration problem. It directly adapts the restoration results at the dense and wide-scale pixel level, requiring accurate discrimination on the rich texture of images. Therefore, in the context of adapting the restoration model to the real-world domain, the diffusion model cannot be extremely simplified.
Figure A9: Failure case of the proposed method and comparison methods.
Figure A10: Analysis of training dynamics (left) and model complexity (right) of the proposed method.
DnCNN Restormer SwinIR
Vanilla Ours Vanilla Ours Vanilla Ours
PSNR 25.98 30.46 24.10 33.98 30.86 34.85
SSIM 0.5911 0.7637 0.5194 0.9183 0.7544 0.9153
Table A3: Performance comparison on SIDD test set of the commonly-used and SOTA restoration models under different training strategies. Vanilla denotes training the model on the paired synthetic datasets and Ours denotes training with the proposed noise-space domain adaptation strategy.
A4.7 Validation on More Restoration Models
Our work contributes to a novel and general domain adaptation strategy for image restoration, which cannot be replaced by current self-supervised methods. To this end, we further validated our method on the commonly used and SOTA restoration models, such as DnCNN, Restormer, and SwinIR. The quantitative evaluations of these comparison methods are reported in Table. As we can observe, all restoration models trained on synthetic datasets failed to generalize well to the real-world dataset. By incorporating the proposed domain adaptation training strategy, the real-world performance of these models gains significant improvements, demonstrating the favorable generalization and scalability of our method.
Algorithm 1 Training Process of The Proposed Framework
0:  Degraded images $\bm{x}^{s}$ and ground truth images ${\bm{y}}^{s}$ from synthetic domain $\mathcal{D}^{s}$, degraded images $\bm{x}^{r}$ from real-world domain $\mathcal{D}^{r}$
0:  Restored synthetic image ${\hat{\bm{y}}}^{s}$ and restored real-world image ${\hat{\bm{y}}}^{r}$
1:  Initialization on the image restoration model $\bm{f}_{\theta}(\cdot)$ and diffusion model $\bm{\epsilon}_{\theta}(\cdot)$
2:  repeat
3:     for all $\bm{x}^{s}_{i}$ and ${\bm{y}}^{s}_{i}$ $\in\mathcal{D}^{s}$, $\bm{x}^{r}_{j}\in\mathcal{D}^{r}$ do
4:        ${\hat{\bm{y}}}^{s}_{i}\leftarrow\bm{f}_{\theta}(\bm{x}^{s}_{i})$, ${\hat{\bm{y}}}^{r}_{j}\leftarrow\bm{f}_{\theta}(\bm{x}^{r}_{j})$
5:        Calculate the restoration loss $\mathcal{L}_{Res}$ with ${\hat{\bm{y}}}^{s}_{i}$ and ${\bm{y}}^{s}_{i}$ using Charbonnier loss
6:        Sample the diffusion’s input $\tilde{\bm{y}}^{s}_{i}=\sqrt{\bar{\alpha}_{t}}{\bm{y}}^{s}_{i}+\sqrt{1-\bar{\alpha}_{t}}\bm{\epsilon}$, $\bm{\epsilon}\sim N(0,\bm{I}),t\in[1,T]$
7:        Predict the noise $\tilde{\bm{\epsilon}}\leftarrow\bm{\epsilon}_{\theta}\left(\tilde{\bm{y}}^{s}_{i}|\mathbf{C}(\hat{\bm{y}}^{s}_{i},\hat{\bm{y}}^{r}_{j}),t\right)$, calculate the diffusion loss $\mathcal{L}_{Dif}$
8:        Update $\bm{\epsilon}_{\theta}(\cdot)$ using $\mathcal{L}_{Dif}$
9:        Update $\bm{f}_{\theta}(\cdot)$ using Eq.4 with $\mathcal{L}_{Res}$ and $\mathcal{L}_{Dif}$
10:     end for
11:  until Convergence
Algorithm 2 One-Pass Inference of The Proposed Framework
0:  Degraded image $\bm{x}^{r}$, trained image restoration model $\bm{f}_{\theta}(\cdot)$
1:  ${\hat{\bm{y}}}^{r}\leftarrow\bm{f}_{\theta}(\bm{x}^{r})$
2:  return  ${\hat{\bm{y}}}^{r}$
Appendix A5 Additional Technological Details
A5.1 Training and Inference
To provide a clear distinction and description of the training and inference stages, we show their detailed processes in Alg. 1 and Alg. 2, respectively. For clarity, we omit the strategies to eliminate the shortcut solutions (exploited in the 7th step of Alg. 1), such as the channel shuffling layer and residual-swapping contrastive learning.
A5.2 Information Guidance between Two Domains
Our motivation for this work is derived from an interesting observation: the prediction error of a conditional diffusion model relies on the quality of the conditions (as shown in Fig. 1 (a)). Therefore, guided by the back-propagated diffusion loss, the restoration network is optimized to provide “good” conditions to minimize the diffusion model’s noise prediction error, aiming for a clean image distribution. During this joint training, the synthetic GT serves as the denoising target in the diffusion model, which potentially offers realistic textures to help adapt the degraded real-world images into the clean distribution. In other words, the clean knowledge/information “leaked” by the diffusion’s input (in a multi-step denoising manner) plays an important role in bridging the gap between different domains.
Generally, the ground truth images and those restored images by a restoration model, whether from synthetic or real-world domains, should reside within a common distribution of high-quality, clean images. However, the appearance of synthetic GT and restored real-world data is unrelated, leading the diffusion model to exploit a shortcut, overfitting its denoising capability by relying solely on the paired synthetic data. To this end, we further design crucial strategies (e.g., channel-shuffling layer and residual-swapping contrastive learning) to implicitly blur the boundaries between conditioned synthetic and real data and prevent the reliance of the model on easily distinguishable features. As a result, the proactive “leakage” from clean distribution to degraded images can effectively work during the whole training process, consistently improving the restoration performance on real-world images.